An air mattress is 2.3 m long, 0.66 m wide and 14 cm deep. if the air mattress itself has a mass of 0.22 kg, what is the maximum mass it can support in freshwater?
Answer #1

Volume V = 2.3 m * 0.66 m * 14 cm = 2.3 m* 0.66 m * 0.14m

                  = 0.21252 m ^ 3

Mass m = 0.22 kg

Let the maximum mass supports be M

Them ( M + m) g = V Dg

           (M+m) = VD

Where D = density of water = 1000 kg / m^ 3

Plug the values we get M + m = 212.52 kg

From this M = 212.3 kg